At Pete’s Pizzeria, a pizza costs $6 and a soda costs 75¢. What is the total cost of 2 pizzas and 4 sodas (before taxes)?
Answer by mananth(16949)

pizza 1-----$6
For 2 it is 2*6 = 12.00
..
soda 1.....$0.75
6 sodas .. 6*0.75= 4.50
..
Total = 12+4.50=$ 16.50